# Benchmarking Unified Face Attack Detection via Hierarchical Prompt Tuning

### v1

@@ II-B Prompt Learning
TABLE I: Comparison of existing UAD datasets. UniAttackData+ surpasses all other datasets in scale and diversity. ”V” and ”I” indicate video- and image-based counts, respectively. The numbers following each attack type in the ”# Catagories” row represent the number of algorithms included in that category.
Dataset # Attack Types # Videos
CASIA-SURF phy.6 21,000
CeFA phy.3 23,538
HiFiMask phy.5 54,600
SiW-v2 phy.14 1,700
MSU-MFSD phy.2 440
CASIA-FASD phy.3 600
ReplayAttack phy.2 1200
OULU-NPU phy.2 5940
FaceForensics++ digi.5 6,000
DeeperForensics-1.0 digi.1 60,000
DFDC digi.8 128,064
ForgeryNet digi.15 221,247
UniAttackData+ phy.10, digi.40 697,347
TABLE II: Comparison of existing PAD and FFD datasets. The scale of our UniAttackData+ is far more than the types in the existing PAD and FFD datasets. The ”#” represents ”the number of”.
III UniAttackDataPlus
III-A Advantages of UniAttackData+
@@
Our UniAttackData+ dataset addresses five critical issues in current UAD datasets, i.e., insufficient sample size, limited attack coverage, and lack of interpretability, providing a comprehensive and valuable benchmark for the UAD community.
@@
Scale. UniAttackData+ includes $2,875$ identities spanning three ethnicities, totaling $697,347$ video segments, with at least $25$ frames per video. Of these, $18,250$ are authentic face videos, 55,950 are physical attack videos, and $623,147$ are digital attack videos. Detailed data is presented in Tab. I and Tab. II. Our UniAttackData+ shows its superiority in separate PAD and FFD datasets, and also in the existing UAD datasets. The large dataset ensures a comprehensive and complete learning space, guaranteeing the effectiveness of detection models from a data perspective.
@@
Attack Types. UniAttackData+ comprises $54$ attack types in total, including $14$ physical attacks and $40$ digital attacks, covering all known attack types in the field. This diversity ensures the model does not overfit to a specific attack type, thereby enhancing its generalization ability. All the attack methods included in our dataset were proposed after 2021, incorporating the latest digital generation attacks. UniAttackData+ ensures the data’s applicability in real-world scenarios, significantly expanding the range of attack types for model training and reducing vulnerability to emerging advanced attacks.
@@
Interpretability. Existing UAD datasets merely merge different datasets without any correlation between the data, attack methods, and identities. This simplistic combination hinders the model from focusing on genuine forgery cues and greatly reduces model robustness in real-world applications. In contrast, UniAttackData+ preserves the identity consistency from UniAttackData while structuring all attack types hierarchically. Each attack method and sample is categorized from coarse to fine and labeled accordingly. This construction method makes full use of prior knowledge, offering researchers an actionable perspective to address UAD challenges. The hierarchical label collection and setup provide a more comprehensive experimental and analytical space, enabling deeper and more interpretable analyses of detection models and algorithms.
III-B Collection of UniAttackData+
@@
Live Face source. The live faces in the UniAttackData+, enhancing diversity in terms of face identity, angle, expression, and ethnicity, are sourced from three public datasets: CASIA-SURF [12], which includes $1,000$ subjects; CASIA-SURF CeFA [13] (CeFA), comprising 1,800 subjects from three ethnicities - African, East Asian, and Central Asian, with $600$ subjects each ethnicity; and CASIA-SURF HiFiMask [14] (HiFiMask), which contains $75$ subjects. These varied sources significantly enhance the dataset’s diversity across multiple dimensions. We merge three existing datasets for the live face data to obtain $2,875$ distinct identities. For CASIA-SURF, the Intel RealSense SR300 camera captures collectors turning left or right, moving up or down, and approaching or moving away from the camera. The same collection method is used for the CASIA-SURF CeFA dataset. HiFiMask contains multiple lighting conditions, such as white, green, and periodic three-colors. Additionally, supplementary light sources from six directions are used to simulate real-world application scenarios.
@@
For dataset preprocessing, we apply MTCNN [15] for face detection and cropping on the real face data from the original datasets, with the scale parameter set to 1.3. We then manually assign class labels, attack type labels, and hierarchical labels to facilitate research.
III-C Spoofing Approach
@@
To ensure diversity in physical attack methods within the UniAttackData+ dataset, we incorporate $14$ physical face spoofing techniques, including $11$ 2D presentation attacks and $3$ 3D mask attacks. The 2D attacks consist of $6$ paper-cutout variations and $2$ print attack types, as well as $3$ video-replay attacks to enrich the dataset diversity further. For 3D mask attacks, UniAttackData+ includes high-fidelity masks made from three different materials: transparent, plaster, and resin. These masks are recorded under varied lighting conditions and with multiple imaging devices for added variation.
@@
Category Type Methods # Videos Total
2D Print Cloth-Indoor 1,800 3,600 55,950
Cloth-Outdoor 1,800
Replay pic-phone 1,800 5,400
screen-phone 1,800
screen-rs 1,800
Cutouts Cutout1 1,000 6,000
Cutout2 1,000
Cutout3 1,000
Cutout4 1,000
Cutout5 1,000
Cutout6 1,000
3D Transparent 13,650 40,950
Plaster 13,650
Resin 13,650
TABLE III: The detailed statistics of physical attacks in UniAttackData+. The ”#” represents ”the number of”.
III-D Forgery Approach
III-D1 Digital Manipulation Attacks
Digital Manipulation Attacks in UniAttackData+ encompasses $17$ specific methods grouped into three types of attacks: (1) Face Swapping [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], where identity features of a source face, $I_{\text{source}}$, are transferred into a target face, $I_{\text{target}}$, producing a composite image, $I_{\text{composite}}=f(I_{\text{source}},I_{\text{target}})$. (2) Facial Attribute Editing [27, 28], which alters specific attributes of a face, such as age, expression, or gender, expressed as $I^{\prime}=g(I,A_{\text{target}})$, where $A_{\text{target}}$ denotes the desired attribute. (3) Video-Driven Facial Animation [29, 30], which generates facial movements in static images using motion parameters $M$ from a driving source, represented by $I_{\text{animated}}=h(I,M)$.
Category Type Methods # Videos Total
Digital Manipulation Attribute-edit SAFA [27] 15,998 32,659 242,859
StarGAN [28] 16,661
Face-Swap FaceDancer [16] 16,168 176,878
GHOST [17] 13,135
SimSwap [18] 10,563
InsightFace [19] 13,264
BlendFace [20] 9,645
DeepLiveCam [31] 12,059
REFace [21] 12,756
FOMM [22] 16,661
InfoSwap [23] 16,534
MegaFS [24] 14,334
HifiFace [25] 16,661
DiffSwap [26] 13,761
Roop [32] 11,337
Video-Driven DaGAN [29] 16,661 33,322
OneShotTH [30] 16,661
TABLE IV: The detailed statistics of digital manipulation attacks.
@@ III-D1 Digital Manipulation Attacks
III-D2 Digital Adversarial Attacks
@@
Category Type Methods # Videos Total
Digital Adversarial Pixel_level ALMA [33] 16,661 183,271 266,576
AdvDrop [34] 16,661
AI_FGTM [35] 16,661
ILA-DA [36] 16,661
AutoAttack [37] 16,661
PGD [38] 16,661
Jitter [40] 16,661
NIFGSM [41] 16,661
PIFGSMpp [42] 16,661
Pixel [43] 16,661
VMIFGSM [44] 16,661
Semantic-Level Demiguise [45] 16,661 83,305
SSAH [46] 16,661
DeCoWA [39] 16,661
PNA_PatchOut [47] 16,661
TGR [48] 16,661
TABLE V: The detailed statistics of digital adversarial attacks.
Pixel-Level. ALMA [33] utilizes Augmented Lagrangian-based attack for efficient, minimally perturbed adversarial examples, combining generality and speed. AdvDrop [34] introduces a novel adversarial attack by dropping imperceptible details from images instead of adding disturbances, challenging DNN robustness and evading existing defense mechanisms. AI-FGTM [35] improves adversarial example generation with dynamic step size, smaller kernels, and a tanh-based gradient method, achieving higher transferability, indistinguishability, and attack success rates than state-of-the-art methods. ILA-DA [36] enhances adversarial example transferability by combining automated image transformations, reverse adversarial updates, and attack interpolation, significantly outperforming prior methods in both defended and undefended scenarios. AutoAttack [37] introduces an ensemble of parameter-free adversarial attacks, improving PGD with optimized step size and objective function to rigorously evaluate adversarial robustness, uncovering weaknesses in many published defenses. PGD [38] explores adversarial robustness through robust optimization, introducing methods for training and attacking neural networks with universal security guarantees against first-order adversaries, significantly enhancing resistance to adversarial attacks. Jitter [40] proposes a novel loss function that improves attack efficiency and success rates by better exploring the perturbation space and addressing model confidence issues. NI-FGSM [41] leverages Nesterov accelerated gradients, and SIM, optimizing perturbations across scaled copies, to enhance the transferability of adversarial examples, achieving higher success rates on defense models. PIM++ [42] enhances adversarial transferability by using a patch-wise iterative method with an amplified step size, a projection kernel for gradient redistribution, and temperature adjustments to balance transferability and white-box attack performance. Pixel [43] introduces a black-box attack that rearranges a small number of pixels to achieve high attack success rates across models and datasets, requiring few iterations and minimal perceptual changes. VMI-FGSM [44] enhances gradient-based adversarial attacks by stabilizing update directions with gradient variance, significantly improving transferability.
@@
III-D3 Digital Generation Attacks
UniAttackData+ comprise $7$ types leveraging diffusion models to produce highly realistic synthetic content. These attacks are divided into three categories: Identity-Consistent Generation, Style Transfer, and Prompt-Driven Generation. Identity-Consistent Generation [49, 50, 51] seeks to maintain a subject’s identity across various contexts, modeled as $I_{\text{gen}}=G(I_{\text{id}},z)$, where $I_{\text{id}}$ is the identity input and $z$ represents variations such as pose or expression. Style Transfer, as in PhotoMaker [52], adapts identity features to different styles, represented as $I_{\text{styled}}=G(I_{\text{id}},\mathbf{S}_{\text{style}})$, to create visually diverse outputs. Prompt-Driven Generation [53] utilizes textual or image prompts to guide synthesis, expressed as $I_{\text{prompted}}=G(P_{\text{text}},P_{\text{img}})$.
ID-Consistant. ConsistentID [49] enhances personalized portrait generation with a multimodal facial prompt generator for fine-grained control and an ID-preservation network with facial attention localization, achieving superior ID consistency and precision from a single reference image. PuLID [50] introduces a tuning-free ID customization method for text-to-image generation, combining a Lightning T2I branch with contrastive alignment and ID losses to ensure high ID fidelity and minimal disruption to original image elements. InstantID [51] introduces a plug-and-play module with IdentityNet, leveraging strong semantic and weak spatial conditions to achieve high-fidelity, single-image-based identity personalization, seamlessly integrating with pre-trained diffusion models like SD1.5 and SDXL.
@@ III-D3 Digital Generation Attacks
Category Type Methods # Videos Total
Digital Generation ID-Consistent ConsistentID [49] 15,436 46,321 113,712
PuLID [50] 16,210
InstantID [51] 14,675
Style Transfer PhotoMaker [52] 18,192 18,192
Prompt Based IP-Adapter_kolors [53] 16,449 49,199
IP-Adapter_sdplusface [53] 16,449
IP-Adapter_sdpxlface [53] 16,301
TABLE VI: The detailed statistics of digital generation attacks.
III-E Protocol
@@
ID Partition Protocol $P1$. Simply merging existing PAD and FFD datasets can cause models to focus on ID noise rather than discriminative features. To address this, we leverage the ID consistency of UniAttackData+. In this protocol, we train the model on a selected subset of identities, using the remaining identities for validation and testing. Each ID contains all attack types.
@@
Cross-attack methods Protocol $P2$. In real-world deployment scenarios, a myriad of attack methods may emerge. Therefore, models should not only fit to seen attack methods but must also be capable of handling unseen attack methods. This protocol is designed to evaluate the model’s generalization ability across different attack methods. We adopted all identities for training and randomly selected 50% of attack methods as the training set, 20% as the validation set, and 30% as the test set. For instance, in the case of digital attacks, we used the first seven types of digital editing face swap attacks as the training set, two types for validation, and four types for testing, in order to assess the model’s performance on unseen face swap attacks.
@@
Cross-attack types Protocol $P3$. In many cases, models are trained only on simpler attack types but are tested against more advanced and complex attack methods. This protocol aims to assess the model’s generalization ability from simpler to more sophisticated attacks. We trained the model using all identities, with 70% of physical 2D attacks, digital editing, and adversarial attacks from each identity as the training set, and 30% as the validation set. Physical 3D attacks and digital generative attacks were used as the test set, to evaluate the model’s generalization ability on unseen, more advanced attack types.
IV Hierarchical Prompt Tuning (HiPTune)

### v3

@@ 2.2 Face Attack Detection Methods
3 UniAttackDataPlus
3.1 Advantages of UniAttackData+
@@
Our UniAttackData+ addresses three critical issues in current UAD datasets, i.e., insufficient sample size, limited attack coverage, and lack of interpretability, providing a comprehensive and valuable benchmark for the UAD community.
@@
Scale. UniAttackData+ includes $2,875$ identities spanning three ethnicities, totaling $697,347$ video segments, with at least $25$ frames per video. Of these, $18,250$ are authentic face videos, 55,950 are physical attack videos, and $623,147$ are digital attack videos. Detailed data is presented in Tab. II and Tab. I. Our UniAttackData+ shows its superiority in separate PAD and FFD datasets, and also in the existing UAD datasets. It ensures a comprehensive and complete learning space, guaranteeing the effectiveness of detection models from a data perspective.
@@
Attack Types. UniAttackData+ comprises $54$ attack types in total, including $14$ physical attacks and $40$ digital attacks, covering all known attack types in the field. This diversity ensures the model does not overfit to a specific attack type, thereby enhancing its generalization ability. All the attack methods included in our dataset were proposed after 2021, incorporating the latest digital generation attacks. UniAttackData+ promotes real-world generalizability by significantly expanding attack diversity, enabling more comprehensive training and reducing model susceptibility to novel advanced attacks.
@@
Interpretability. Existing UAD datasets merely merge different datasets without any correlation between the data, attack methods, and identities. This simplistic combination hinders the model from focusing on genuine forgery cues and greatly reduces model robustness in real-world applications. Unlike previous datasets, UniAttackData+ maintains the identity consistency of UniAttackData [39] and further organizes all attack types into a hierarchical taxonomy, enabling each sample to be labeled from coarse to fine levels. The hierarchical labeling scheme offers a richer experimental and analytical framework, facilitating more interpretable and fine-grained evaluation of detection models.
3.2 Collection of UniAttackData+
@@
Live Face source. The live faces in the UniAttackData+, enhancing diversity in terms of face identity, angle, expression, and ethnicity, are sourced from three public datasets: CASIA-SURF [1], which includes $1,000$ subjects; CASIA-SURF CeFA [56] (CeFA), comprising 1,800 subjects from three ethnicities, i.e., African, East Asian, and Central Asian, with $600$ subjects each ethnicity; and CASIA-SURF HiFiMask [3] (HiFiMask), which contains $75$ subjects. These varied sources significantly enhance the dataset’s diversity across multiple dimensions. We merge three existing datasets for the live face data to obtain $2,875$ distinct identities. CASIA-SURF and CeFA use the RealSense SR300 to capture head movements and distance variations. HiFiMask simulates real-world lighting with white, green, and tri-color cycles, plus six-directional supplementary lights.
@@
Preprocessing. For dataset preprocessing, we apply MTCNN [60] for face detection and cropping on the real face data from the original datasets, with the scale parameter set to 1.3. We then manually assign class labels, attack type labels, and hierarchical labels to facilitate research.
3.3 Spoofing Approach
@@
To ensure diversity in physical attack methods within the UniAttackData+ dataset, we incorporate $14$ physical face spoofing techniques, including $11$ 2D presentation attacks and $3$ 3D mask attacks. See from the Tab. 3, The 2D attacks consist of $6$ paper-cutout variations and $2$ print attack types, as well as $3$ video-replay attacks. For 3D mask attacks, UniAttackData+ includes high-fidelity masks made from $3$ different materials: transparent, plaster, and resin. These masks are recorded under varied lighting conditions and with multiple imaging devices for added variation.
@@
Figure 3: This diagram categorizes attacks into four distinct types: Physical Attacks, Digital Manipulation, Digital Adversarial and Digital Generation. Physical Attacks comprise 14 subtypes, Digital Manipulation includes 17 subtypes, the Digital Adversarial group contains 16 subtypes, and the Digital Generation group has 7 subtypes. We can clearly see two advantages of datasets containing forgery methods: (2) rich and diverse types of attacks; (2) Determine the proportion based on the commonness and progressiveness of attack types.
3.4 Forgery Approach
3.4.1 Digital Manipulation Attacks
As shown in Tab. 3, Digital Manipulation Attacks in UniAttackData+ encompasses $17$ specific methods grouped into three types of attacks: (1) Face Swapping [20, 61, 21, 62, 63, 64, 65, 66, 67, 68, 69], where identity features of a source face, $I_{\text{source}}$, are transferred into a target face, $I_{\text{target}}$, producing a composite image, $I_{\text{composite}}=f(I_{\text{source}},I_{\text{target}})$. (2) Facial Attribute Editing [18, 19], which alters specific attributes of a face, such as age, expression, or gender, expressed as $I^{\prime}=g(I,A_{\text{target}})$, where $A_{\text{target}}$ denotes the desired attribute. (3) Video-Driven Facial Animation [22, 23], which generates facial movements in static images using motion parameters $M$ from a driving source, represented by $I_{\text{animated}}=h(I,M)$.
@@ 3.4.1 Digital Manipulation Attacks
TABLE III: Sample statistics for each protocol on the UniAttackData+ dataset, including train, dev, and test sets, where each column represents the number of samples for live faces or fake types. ‘#’ indicates the number of videos.
Protocol Subset #Live Physical Attacks Digtal Manipulation Digital Adversarial Digital Generation #All
#2D #3D
#Attribute-
Edit
#Face-
Swap
#Video-
Driven
#Pixel-
Level
#Semantic-
Level
#ID-
Consistent
#Style-
Transfer
#Prompt-
Based
P1 train 9,125 7,500 20,475 12,330 88,439 16,661 91,636 41,653 23,161 9,096 24,600 355,676
dev 3,650 3,000 8,190 6,532 35,376 6,664 36,654 16,661 9,264 3,638 9,840 139,469
text 5,475 4,500 12,285 13,797 53,063 9,997 54,981 24,991 13,896 5,458 14,759 213,202
P2 train 13,650 5,880 19,110 11,663 89,522 11,663 81,639 34,988 21,078 9,096 23,026 321,315
dev 1,000 2,520 8,190 4,998 38,366 4,998 34,988 14,995 9,033 3,638 9,868 132,594
text 3,600 6,600 13,650 15,998 48,990 16,661 49,983 33,322 16,210 5,458 16,301 226,773
P3 train 13,650 10,500 - 22,861 123,815 23,325 128,290 58,314 - - - 380,755
dev 1,000 4,500 - 9,798 53,063 9,997 54,981 24,991 - - - 158,330
text 3,600 - 40,950 - - - - - 46,321 18,192 49,199 158,262
3.4.2 Digital Adversarial Attacks
@@
Pixel-Level. ALMA [72] utilizes Augmented Lagrangian-based attack for efficient, minimally perturbed adversarial examples, combining generality and speed. AdvDrop [73] introduces a novel adversarial attack by dropping imperceptible details from images instead of adding disturbances, challenging DNN robustness and evading existing defense mechanisms. AI-FGTM [24] improves adversarial example generation with dynamic step size, smaller kernels, and a tanh-based gradient method, achieving higher transferability, indistinguishability, and attack success rates than state-of-the-art methods. ILA-DA [25] enhances adversarial example transferability by combining automated image transformations, reverse adversarial updates, and attack interpolation, significantly outperforming prior methods in both defended and undefended scenarios. AutoAttack [74] introduces an ensemble of parameter-free adversarial attacks, improving PGD with optimized step size and objective function to rigorously evaluate adversarial robustness, uncovering weaknesses in many published defenses. PGD [75] explores adversarial robustness through robust optimization, introducing methods for training and attacking neural networks with universal security guarantees against first-order adversaries, significantly enhancing resistance to adversarial attacks. Jitter [27] proposes a novel loss function that improves attack efficiency and success rates by better exploring the perturbation space and addressing model confidence issues. NI-FGSM [76] leverages Nesterov accelerated gradients and SIM, optimizing perturbations across scaled copies, to enhance the transferability of adversarial examples, achieving higher success rates on defense models. PIM++ [77] enhances adversarial transferability by using a patch-wise iterative method with an amplified step size, a projection kernel for gradient redistribution, and temperature adjustments to balance transferability and white-box attack performance. Pixel [78] introduces a black-box attack that rearranges a small number of pixels to achieve high attack success rates across models and datasets, requiring few iterations and minimal perceptual changes. VMI-FGSM [79] enhances gradient-based adversarial attacks by stabilizing update directions with gradient variance, significantly improving transferability.
@@
3.4.3 Digital Generation Attacks
UniAttackData+ comprises $7$ types leveraging diffusion models to produce highly realistic synthetic content. These attacks are divided into three categories: Identity-Consistent Generation, Style Transfer, and Prompt-Driven Generation. Identity-Consistent Generation [31, 32, 33] seeks to maintain a subject’s identity across various contexts, modeled as $I_{\text{gen}}=G(I_{\text{id}},z)$, where $I_{\text{id}}$ is the identity input and $z$ represents variations such as pose or expression. Style Transfer, as in PhotoMaker [81], adapts identity features to different styles, represented as $I_{\text{styled}}=G(I_{\text{id}},\mathbf{S}_{\text{style}})$, to create visually diverse outputs. Prompt-Driven Generation [82] utilizes textual or image prompts to guide synthesis, expressed as $I_{\text{prompted}}=G(P_{\text{text}},P_{\text{img}})$.
ID-Consistent. ConsistentID [31] enhances personalized portrait generation with a multimodal facial prompt generator for fine-grained control and an ID-preservation network with facial attention localization, achieving superior ID consistency and precision from a single reference image. PuLID [32] introduces a tuning-free ID customization method for text-to-image generation, combining a Lightning T2I branch with contrastive alignment and ID losses to ensure high ID fidelity and minimal disruption to original image elements. InstantID [33] introduces a plug-and-play module with IdentityNet, leveraging strong semantic and weak spatial conditions to achieve high-fidelity, single-image-based identity personalization, seamlessly integrating with pre-trained diffusion models like SD1.5 and SDXL.
@@ 3.4.3 Digital Generation Attacks
3.5 Protocol and Statistic
@@
Identity Division Protocol $P1$. Simply merging existing PAD and FFD datasets can lead models to focus on identity-related noise rather than truly discriminative features. To mitigate this issue, we leverage the ID consistency provided by UniAttackData+. In this protocol, the model is trained on a selected subset of identities, while the remaining identities are reserved for validation and testing. Each identity includes all types of attacks to ensure comprehensive evaluation.
@@
Cross-attack Methods Protocol $P2$. In real-world deployment scenarios, a wide range of attack methods may emerge. Therefore, models must not only perform well on known attack types but also generalize effectively to unseen ones. This protocol is designed to evaluate a model’s generalization capability across diverse attack methods. All identities are used for training, while attack methods are split randomly: 50% for training, 20% for validation, and 30% for testing. For example, in the case of digital attacks, the first seven types of digital face, i.e., swap attacks, are used for training, two types for validation, and four types for testing, enabling us to assess the model’s ability to handle unseen face-swap attacks.
@@
Cross-attack types Protocol $P3$. In many cases, models are trained only on simpler attack types but are tested against more advanced and complex attack types. This protocol aims to assess the model’s generalization ability from simpler to more sophisticated attacks. We trained the model using all identities, with 70% of physical 2D attacks, digital editing, and adversarial attacks from each identity as the training set, and 30% as the validation set. Physical 3D attacks and digital generative attacks were used as the test set to evaluate the model’s generalization ability on unseen, more advanced attack types.
The sample statistics for each protocol are shown in Tab. III. We have detailed the number of train, dev, and test sets for each protocol, and also listed the types of spoofing contained in the subsets and their quantities.
4 Hierarchical Prompt Tuning (HiPTune)